TRYING TO SOLVE THIS PROBLEM
-4(a+3)-a=12

Answer by dlh(17)

-4(a+3)-a=12 Original problem
-4a-12-a=12 Distributive property
-5a-12=12 Simplify
-5a=24 Adding to other side
a=24/-4 Dividing to other side

I added an equal sign because I wouldn't be able to solve it other wise.